OpenStudy (firejay5):

Find the LCM of each set of polynomials. If needed, show work and explain. 18. 4w - 12, 2w - 6

OpenStudy (anonymous):

\[ LCD(a,b) = \frac{a\times b}{GCD(a,b)} \]

OpenStudy (anonymous):

Since \(4w-12 = (2)(2w-6)\), we know the GCF is \(2w-6\) \[\begin{split} LCM(4w-12,2w-6) &= \frac{(4w-12)(2w-6)}{GCF(4w-12,2w-6)}\\ &= \frac{(4w-12)(2w-6)}{2w-6} \\&= 4w-12 \end{split}\]

OpenStudy (firejay5):

What's the answer @satellite73 and @Mertsj

OpenStudy (mertsj):

Factor each one and take the largest power of every DIFFERENT factor. That is the LCM

OpenStudy (mertsj):